OpenStudy (anonymous):

cos(285[degrees]) Exact value?

OpenStudy (precal):

just put it in the calculator? .2588190451

OpenStudy (agent0smith):

precal they need an exact value. cos285 = cos75 Now you can use angle sum formulas, since cos75 = cos(45+30) cos(A+B) = cosAcosB + sinAsinB cos(30+45) = cos30cos45 + sin30sin45 And you can use an exact values chart for those and solve it. This should help if you get stuck: http://www.mathwarehouse.com/trigonometry/identities/angle-sum-formula.php

OpenStudy (precal):



OpenStudy (precal):

Even/Odd Identities cos(-75)=cos75

OpenStudy (precal):